OpenStudy (anonymous):

At his trucking company, Alfred received an order to ship 297 cubic feet of oil in cylindrical steel barrels. If each barrel has a radius of 1.2 feet and is 5 feet tall, how many steel barrels will Alfred need? 12 13 14 15

OpenStudy (anonymous):

The exercise did not provide the value of pi ?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

Good is 297 m ^ 3 which will be distributed in cylindrical barrels that have 1.2 radius 5m tall!   ha logic is simple you think the volume of barrels then divide the total amount of oil volume by the barrel. looks like this: V = ^ π.r 2.h V = (3,14). (1.2 ^ 2) (5). V = 22.608 297 / V = amount of barrel 297/22, 608 = approximately 13

OpenStudy (anonymous):

thanks so much c: